A family vacation trip took 12 hours. Of the total, 4/5 was in car and 1/5 was in boat, and the car traveled 5 times as fast as the boat. How much more time, in hours, was spend in the boat than the car?
Found 2 solutions by ikleyn, richwmiller:
Answer by ikleyn(52855)

.
The condition is presented in so ambiguous way that . . .

I can not understand if 4/5 is the fraction of time or the fraction of the distance.

Sorry, it is not the Math way to formulate the problem in this manner.


Answer by richwmiller(17219)

Grammatically, the problem states that 4/5 and 1/5 are time.
However, the question itself indicates that more time was spent in the boat which contradicts that premise.
So, the problem contradicts itself.
4/5*12=48/5 =9 3/5 hrs=9:36 hrs time in car
1/5*12=12/5= 2 2/5 hrs=2:24 hrs time in boat
9:36-2:24= 7:12 hrs more in the car than in the boat
Another way to calculate:
4/5-1/5=3/5
3/5*12=36/5= 1/5= 7:12 hrs more in the car than in the boat